OpenStudy (anonymous):

art 1: Show the work to solve |x+3| +7 10 (3 points) Part 2: Describe the graph of the solution in a complete sentence. (3 points)

OpenStudy (anonymous):

Please answer !! <3

OpenStudy (lgbasallote):

can you reWRITE the question? especially that "art 1"

OpenStudy (anonymous):

Show the work to solve |x+3|+7\[|x+3|+7\le10\]

OpenStudy (lgbasallote):

ahh better first subtract 7 from both sides \[|x+3| \le 3\] now solve for x when \[x + 3 \le 3\] AND \[x +3 \ge -3\]

OpenStudy (lgbasallote):

when \[x + 3 \le 3\] then simply subtract 3 from both sides this gives \[x \le 3 - 3\] simplify that ;)

OpenStudy (lgbasallote):

when \[x + 3 \ge -3\] subtract 3 from both sides again \[x \ge -3 -3\] simplify that too ;)

OpenStudy (anonymous):

I got this part thx , i need help answering part 2 ....Thx 4 helping me

OpenStudy (lgbasallote):

well you can only do this when you have a graph...what are the x values you got?

OpenStudy (lgbasallote):

by that i mean \[x \le ?\] what did you get for this \[x \ge ?\] what did you get for this

OpenStudy (anonymous):

\[x \ge0\]

OpenStudy (anonymous):

\[x \le6\]

OpenStudy (lgbasallote):

hmm...actually it's supposed to be \(x \le 0\) \[x \le 3 - 3\] \[x \le 0\] and it should be \[x \ge -6\] because \[x \ge -3 - 3\] \[x \ge -6\] got it?

OpenStudy (anonymous):

ohh ok thx i keep doing the inverse .. thx :)

OpenStudy (lgbasallote):

so the graph for that is \(-6 \le x \le 0\) therefore...  now describe it